golban:

Susie has 8 over 9cup of ice cream. How many 1 over 4cup servings can she make? 2 over 9 9 over 32 3 and 5 over 9 1 and 8 over 15

golban:

ok

justus:

(8/9) / (1/4) = 8/9 * 4/1 = 32 / 9 =

justus:

Solve this to get the answer 32 / 9 =

justus:

??

justus:

You got it